Trace the curve y^2=(x+1)(x-1)^2 by showing all properties
1
Expert's answer
2019-09-16T10:22:11-0400

The curve is "y^2=(x+1)(x-1)^2"

"(x+1)(x-1)^2\geq 0 \Rightarrow (x+1)\geq 0 \Rightarrow x\geq -1 ."


"y=\pm \sqrt{(x+1)(x-1)^2}"


The graph of the curve is symmetric about the x-axis.


Graph of function "y= \sqrt{(x+1)(x-1)^2}" :

"y(-1)= 0, y(-1)= 0."


"y'(x)= \frac{3x^2+2x-1}{ \sqrt{(x+1)(x-1)^2}}"

"y'(-1\/3)= 0 ;"


"y'(1)" - not exist.


The function increases for x in the interval "[-1;-\frac{1}{3}] and [1;\infty], (y'>0)"


The function decreasing for x in the interval "[-\frac{1}{3};1], (y'<0)"


Let us try to find where a function is increasing or decreasing









Graph of the curve :